Consider a system that has a translation look-aside buffer (TLB) that has a hit-ratio of 85%. On a hit, it only takes 75 nanoseconds to access the memory, however, on a miss, it takes 200 nanoseconds to access the memory. Given the system described above, what is the effective access time?
Answer #1

if the required page address is present in the TLB then take that physical addess then goto main memory and get that word in case if it is not present then we need to goto main memory with some miss rate

then EMAT= t + m + p(k*m)

where t: time taken to access to TLB

m: main memory access time

k: number of page levels

p: miss rate

given p=(1-0.85)=0.15

so EMAT=75+200+0.15(1*200)=305nsec.